### v1

# Discrete time crystal for periodic-field sensing with quantum-enhanced precision

Abstract
Sensing periodic-fields using quantum sensors has been an active field of research. In many of these scenarios, the quantum state of the probe is flipped regularly by the application of $\pi$-pulses to accumulate information about the target periodic-field. The emergence of a discrete time crystalline phase, as a nonequilibrium phase of matter, naturally provides oscillations in a many-body system with an inherent controllable frequency. They benefit from long coherence time and robustness against imperfections, which makes them excellent potential quantum sensors. In this paper, through theoretical and numerical analysis, we show that a disorder-free discrete time crystal probe can reach the ultimate achievable precision for sensing a periodic-field. As the amplitude of the periodic-field increases, the discrete time crystalline order diminishes, and the performance of the probe decreases remarkably. Nevertheless, the obtained quantum enhancement in the discrete time crystal phase shows robustness against different imperfections in the protocol. Finally, we propose the implementation of our protocol in ultra-cold atoms in optical lattices.
I Introduction
Metrological tasks can harness quantum features, such as superposition and entanglement, to achieve greater precision compared to their classical counterparts [1]. To estimate an unknown parameter $h_{a}$, the precision of the estimation $\delta h_{a}$, quantified by standard deviation, is fundamentally lower bounded by Cramér-Rao inequality [2, 3] as $\delta h_{a}{\geq}1/\sqrt{M\mathcal{F}_{Q}}$ where $M$ is the number of observations and $\mathcal{F}_{Q}$ is the quantum Fisher information (QFI) [4]. If the Hamiltonian $H{=}H_{0}{+}h_{a}H_{1}$, in which $H_{0}$ independently describes the physics of the probe and $H_{1}$ is related to the parameter, encodes $h_{a}$ through the unitary evolution $U(t){=}e^{-itH}$, it has been shown that the QFI is upper bounded by $\mathcal{F}_{Q}{\leq}t^{2}||H_{1}||^{2}$, where $||H_{1}||$ is the spectral norm of $H_{1}$ and can be related to the size of the probe $L$ as $||H_{1}||^{2}{\propto}L^{\beta}$ [5, 6, 7]. Although in the absence of quantum features, one obtains $\beta{=}1$, known as the standard quantum limit, their presence allows quantum-enhanced sensitivity, namely $\beta{>}1$. For $H_{0}{=}0$. This quantum enhancement can be achieved in phase shift detection by initializing the probe in an entangled Greenberger–Horne–Zeilinger (GHZ) state [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. However, this approach is limited by the difficulty in creating and maintaining such GHZ-type states, especially on a large scale [18, 19]. In or out of equilibrium, Strongly correlated many-body systems provide a valid alternative to this approach [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. Due to $H_{0}{\neq}0$, many-body sensors benefit from the entanglement that is either intrinsically present in the spectrum or generated freely during the time evolution  [35].
Quantum sensors have also been developed for periodic field sensing inferring the amplitude [39, 40, 41], the frequency [42, 43] and the phase [44, 45, 46, 47, 48, 49] of oscillatory fields in the environment. In many such schemes, high-frequency spin echo pulses, typically used in dynamic decoupling [50, 51, 51, 52, 43, 53, 1] to overcome decoherence, are used to improve the accumulation of information about the time-varying signal in the quantum state of the probe. However, these high-frequency pulses often diminish spin correlations in the quantum state of the probe, which may decrease its sensing capability. Recent studies demonstrate that the spin correlations created by the system’s dynamics can be utilized to achieve enhanced periodic-field sensing [54, 55]. Among these proposals, systems with long-range spatiotemporal ordering, known as time crystals, are attracting increasing interest [56, 57, 58, 59, 60, 61].
Breaking time-translational symmetry in periodically driven non-equilibrium systems may create a new phase of matter, known as time crystalline phase [62, 63, 64, 65, 66, 67, 68]. In recent years, discrete time crystals (DTCs), have been the subject of intensive theoretical [66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93] and experimental [94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110] investigations. In periodically driven systems with a period $T$, DTCs are characterized by subharmonic responses to periodic drives. These responses are robust against small driving imperfections, and are eternal in the thermodynamic limit. In addition to the external drive, one can induce non-equilibrium dynamics through quantum quench to form a DTC in a disorder-free system with Hermitian [72, 111] or even non-Hermitian physics [111]. In closed systems, the existence of DTC order hinges on mechanisms that prevent energy absorption from external driving pulses, such as many-body localization [67, 68, 69, 71, 88, 87, 57], bosonic self-trapping [112, 66, 70, 113, 73, 78, 89], Stark gradient fields [83, 74], gradient interaction [61], domain-wall confinement [79], and quantum scars [76, 82, 81, 80, 84, 106].
A well-known characteristic of DTCs is the formation of paired GHZ-type states with a $\pi$ quasienergy gap in their spectrum. The emergence of these large-scale entangled states that are responsible for subharmonic responses makes DTCs a promising candidate for quantum technological tasks. Thus far, DTCs have found applications in simulating complex systems [75], topologically protected quantum computation [114], quantum engine design [115], and metrology [59, 57, 61, 110]. Sensing in fully connected graphs [59] and gradient interacting systems [61] demonstrates the potential of DTCs for quantum-enhanced precision by achieving $\beta{=}3$ in sensing the exchange coupling While this surpasses the standard quantum limit, the ultimate limit determined by $||H_{1}||^{2}{\propto}L^{\beta}$ with $\beta{=}4$ in sensing the strength of two-body interactions, has not been obtained. Relying on the long coherence time of DTCs, theoretical [57] and experimental [110] results demonstrate the capability of DTCs for constructing highly frequency-selective sensors of AC magnetic fields. Nevertheless, effectively exploiting the inherent GHZ-type eigenstates in the spectrum of the DTCs is a challenge that has yet to be addressed.
In this work, through concrete theoretical and numerical analysis, we show that a devised probe in the DTC order can achieve ultimate estimation precision in a gradient periodic magnetic field sensing. The external periodic-field breaks site invariance symmetry and eliminates the DTC order with a rate that depends on the probe size. However, for small values of the periodic-field, our results demonstrate that in the DTC order, the probe can benefit from the inherent GHZ state in the spectrum and reach ultimate precision. The robustness of this performance has been analyzed against various types of defects and imperfections. We further provide details of our probe’s experimental implementation on ultra-cold atoms in optical lattices.
II Quantum parameter estimation
In this section, we review the theory of quantum parameter estimation. This field seeks to determine an unknown parameter, $h_{a}$, encoded in the quantum state $\rho(h_{a})$. The key objective is to minimize the uncertainty in estimating $h_{a}$, represented by the standard deviation $\delta h_{a}$. This uncertainty is fundamentally limited by the Cramér-Rao inequality [2, 3], which relates the minimum error attainable by any locally unbiased estimator to the classical Fisher information (CFI) $\mathcal{F}_{C}(h_{a})$, as
$\delta h_{a}{\geq}\frac{1}{\sqrt{M\mathcal{F}_{C}(h_{a})}}.$ (1)
where $M$ represents the number of repetitions of the sensing protocol. For a given set of Positive Operator-Valued Measurement (POVM) operators $\{\Pi_{r}\}$, the probability of a measurement outcome $r$ is given by $p_{r}(h_{a}){=}{\rm Tr}[\rho(h_{a})\Pi_{r}]$. One can obtain the CFI as $\mathcal{F}_{C}(h_{a}){=}\sum_{r}p_{r}(h_{a})[\partial_{h_{a}}\ln p_{r}(h_{a})]^{2}$, with $\partial_{h_{a}}$ being the derivatives with respect to the parameter $h_{a}$. Optimizing over all possible POVMs $\{\Pi_{r}\}$ leads to the quantum Cramér-Rao inequality, which is bounded by the quantum Fisher information (QFI) $\mathcal{F}_{Q}(h_{a})$ as
$\delta h_{a}{\geq}\frac{1}{\sqrt{M\mathcal{F}_{C}(h_{a})}}{\geq}\frac{1}{\sqrt{M\mathcal{F}_{Q}(h_{a})}}.$ (2)
Thus, the QFI, as a central quantity in metrology, defines the maximum precision limit in quantum sensing. For pure quantum states $\rho(h_{a}){=}|\psi(h_{a})\rangle\langle\psi(h_{a})|$, the QFI takes a simple form as [4]
$\mathcal{F}_{Q}(h_{a}){=}4\left(\langle\partial_{h_{a}}\psi(h_{a})|\partial_{h_{a}}\psi(h_{a})\rangle{-}|\langle\partial_{h_{a}}\psi(h_{a})|\psi(h_{a})\rangle|^{2}\right).$ (3)
The performance of a sensor is characterized by how the QFI scales with available resources, such as probe size and time. In particular, in closed systems, this relationship can be expressed as $\mathcal{F}_{Q}{\sim}n^{\alpha}L^{\beta}$, where $n$ and $L$ represent the duration time of the protocol and the size of the probe, respectively, with $\alpha$ and $\beta$ as the corresponding exponents. Without quantum effects, the QFI can scale at most quadratically with $n$ and linearly with $L$, represented as $\alpha{=}1$ and $\beta{=}1$, which is known as the standard limit. However, by exploiting quantum probes with quantum features, it is possible to enhance the scaling to $\alpha{>}1$ and $\beta{>}1$, a phenomenon called quantum-enhanced sensitivity. A notable case is when $\alpha{=}2$ and $\beta{=}2$, which is called Heisenberg scaling. The most common way to capture this scaling is by initializing a dynamical protocol in maximally entangled states, such as Greenberger–Horne–Zeilinger (GHZ) states, and encoding the desired parameter through a unitary evolution, while the interaction between the system’s components remains negligible. The limitations of this method prompt efforts to unlock the potential of many-body interacting systems. Strongly correlated many-body systems show a remarkable potential to go beyond the standard limit, with $\beta{>}1$, by harnessing a variety of quantum features such as criticality. Nonetheless, the obtainable precision is fundamentally bounded by the interaction details in the system [5, 6, 7].
III Model
Figure 1: (a) Schematic for the DTC sensor introduced here. (b) Imbalance in the total magnetization of the chains $a$ and $b$ denoted as $\mathcal{I}(nT)$, versus the amplitude of the periodic-field $h_{a}$ at different stroboscopic times in a system of size $L{=}6$ with $\varepsilon{=}0.1$.
Our system consists of two one-dimensional chains, labeled $a$ and $b$, each containing $L$ spin-$1/2$ particles. The system is under binary quench and the Hamiltonian reads
$H(t)=\begin{cases}H_{a}+H_{b}=-J_{z}\sum_{\mu=a}^{b}\sum_{j=1}^{L-1}\sigma_{j}^{\mu z}\sigma_{j+1}^{\mu z}&0<t\leqslant t_{1}\\ H_{I}=J_{ab}\sum_{j=1}^{L}(\sigma_{j}^{a+}\sigma_{j}^{b-}+\sigma_{j}^{a-}\sigma_{j}^{b+})&t_{1}<t\leqslant t_{1}+t_{2}\end{cases}$ (4)
where $J_{z}$ denotes the spin exchange coupling, $\sigma^{\mu\pm}_{j}{=}\frac{1}{2}(\sigma^{\mu x}_{j}{\pm}i\sigma^{\mu y}_{j})$, and $\sigma^{\mu(x,y,z)}_{j}$ representing Pauli operators acting at site $j$ of chain $\mu{=}a,b$. We set $J_{ab}{=}\pi J_{z}(1{-}\varepsilon)$, with $\varepsilon{\in}\mathds{R}$ as a small deviation from a perfect coupling between two chains and $t_{1}{=}t_{2}{=}1{/}2J_{z}$. This binary quench creates a time-crystalline temporal order with periodicity $T_{dtc}{=}2T$ in which $T{=}t_{1}{+}t_{2}$, in a system initialized in $|\psi(0)\rangle{=}|{\uparrow}{\cdots}{\uparrow}\rangle_{a}{\otimes}|{\downarrow}{\cdots}{\downarrow}\rangle_{b}$  [72, 111]. This means that, for infinitesimal $\varepsilon$, measuring the imbalance in the total magnetization of chains $a$ and $b$,
$\mathcal{I}(nT){=}\frac{1}{\mathcal{I}(0)}\sum_{j=1}^{L}\langle\psi(nT)|(\sigma_{j}^{az}{-}\sigma_{j}^{bz})|\psi(nT)\rangle$ (5)
shows persistent coherent oscillation between $\mathcal{I}(nT){=}1$ at even $n$’s obtained for $|\psi(0)\rangle$, and $\mathcal{I}(nT){=}{-}1$ at odd $n$’s obtained for the inversion partner of the initial state $|\tilde{\psi}(0)\rangle{=}\Pi_{j=1}^{L}\sigma^{ax}_{j}{\otimes}\sigma^{bx}_{j}|\psi(0)\rangle$. The origin of this period-doubling oscillations that last for $\tau{\propto}e^{\alpha L}$ with $\alpha{=}\ln(\varepsilon^{-1.033})$ is the emergence of $\pi$-Paired Floquet states as $|\Phi_{\pm}\rangle{\cong}(|\psi(0)\rangle{\pm}|\tilde{\psi}(0)\rangle){/}\sqrt{2}$ with eigenvalues differing by $\pi$ [111].
To investigate the sensing capability of this system, we expose the chain $a$ to a gradient periodic-field in the $z$-direction as
$H_{ac}(t){=}B_{a}(t)G^{az}{=}h_{a}\sin(2\pi f_{ac}t+\theta_{ac})\sum_{j=1}^{L}j\sigma_{j}^{az},$ (6)
where $h_{a}$, $f_{ac}$, and $\theta_{ac}{\in}[0,\pi/2)$ are amplitude, frequency, and the phase of the periodic-field, respectively. See Fig. 1 (a) for a system schematic. The cross-talk effect will be discussed later. The time propagator for one period cycle can be obtained as
[…]
where $\mathcal{T}$ denotes the time-ordering operator. For the periodic-field that resonates with the binary quench, namely $T_{ac}{=}T$, one has $\Theta_{n}{=}\frac{(-1)^{n}h_{a}}{\pi J_{z}}\cos(\theta_{ac})$ for $t_{1}{=}t_{2}{=}1{/}2J_{z}$. Without loss generality, we set $J_{z}{=}1$, and $\theta_{ac}{=}0$. To show the effect of the implemented periodic-field on the DTC order, Fig. 1 (b) plots the imbalance $\mathcal{I}(nT)$ versus the amplitude of the periodic-field $h_{a}$ over tens of stroboscopic times $n$ when the system initialized in $|\psi(0)\rangle$. By increasing $h_{a}$,the system goes through a sharp phase transition from a stable DTC order to a region with no spontaneous breaking of DTTS in Eq. (4). Surprisingly, this transition, happening in a specific value denoted by $h_{a}{=}h_{a}^{\max}$ (dashed line), repeats in all the stroboscopic times, and, as we show latter, is robust against the deviation parameter $\varepsilon$.
In the following, we first show that this phase transition can serve as a valuable resource for quantum-enhanced sensitivity and provide ultraprecise periodic-field sensing. Then, we complete our study by providing a robustness analysis of this enhancement to the perturbations in initialization, frequency offset, and crosstalk. We also explore the effect of $\varepsilon$ on the capability of our probe and propose an experimental protocol for its implementation.
IV DTC sensor
Figure 2: (a) The QFI as a function of $h_{a}$ over different stroboscopic times. (b) The dynamical behavior of the QFI versus stroboscopic time $n$ at different values of $h_{a}$. In both panels (a) and (b), we set $L{=}7$ and $\varepsilon{=}0.1$. (c) QFI as a function of $h_{a}$ for systems of different sizes over $n{=}10$ period cycles when $\varepsilon{=}0.1$. (d) The values of QFI inside the DTC phase, namely $h_{a}{=}10^{-5}$, and in the transition point, namely $h_{a}{=}h_{a}^{\max}$, as a function of the system size $L$. The markers are the numerical results, and the solid lines provide the best fitting function as $\mathcal{F}_{Q}{\propto}L^{\beta}$ with $\beta{=}4.0902$ and $\beta{=}2.5887$ for a sensor in the DTC phase and in its transition point. The inset is the transition point $h_{a}^{\max}$ as a function of $L$. The best fitting function is obtained as $h_{a}^{\max}{\propto}L^{-1.0916}$.
[…]
$\partial_{h_{a}}U_{t}$ $={-i}U_{t}\sum_{k=1}^{n}\Big{(}\Pi_{l=1}^{k}U_{l}\Big{)}^{\dagger}\Big{(}\partial_{h_{a}}\Theta_{k}G^{az}\Big{)}\Big{(}\Pi_{l=1}^{k}U_{l}\Big{)}\leqslant iU_{t}G^{az}/\pi.$ (8)
Here, we used the following simplification
$-i\Big{(}\partial_{h_{a}}$ $\Theta_{k}\Big{)}\Big{(}\Pi_{l=1}^{k}U_{l}\Big{)}^{\dagger}G^{az}\Big{(}\Pi_{l=1}^{k}U_{l}\Big{)}$ (9)
=$=\begin{cases}i(-1)^{k+1}G^{az}/\pi&(k-1)T<t_{k}\leqslant(k-1/2)T\\ i(-1)^{k+1}D_{k}^{az}/\pi&(k-1/2)T<t_{k}\leqslant kT\\ \end{cases}$ (10)
wherein $D_{k}^{az}{=}\sum_{j=1}^{L}j\alpha_{j}^{k}\sigma^{az}_{j}{\leqslant}\sum_{j=1}^{L}j\sigma^{az}_{j}{=}G^{az}$ and $\alpha_{j}{=}\Big{(}j^{2}\Theta_{k}^{2}+J_{ab}^{2}\cos^{2}(\sqrt{j^{2}\Theta_{k}^{2}+J_{ab}^{2}})\Big{)}/\Big{(}j^{2}\Theta_{k}^{2}+J_{ab}^{2}\Big{)}$. Using
$\mathcal{F}_{Q}(h_{a})$ $=$ (11)
44 44 $\left(\langle\psi(0)|(\partial_{h_{a}}U_{t})^{\dagger}(\partial_{h_{a}}U_{t})|\psi(0)\rangle{-}|\langle\psi(0)|U_{t}^{\dagger}(\partial_{h_{a}}U_{t})|\psi(0)\rangle|^{2}\right)$
and Eq. 8, one can simply obtain $\mathcal{F}_{Q}(t){\leqslant}4n^{2}\Delta G^{az}$ in which $\Delta G^{az}{=}\sqrt{\langle(G^{az})^{2}\rangle{-}\langle G^{az}\rangle^{2}}$ stands for the variance that is upper bounded by the square of the spectral norm defined as $||G^{az}||^{2}{=}(\lambda_{\max}{-}\lambda_{\min})^{2}$ with $\lambda_{\max}{=}L(L{-}1)/2$ and $\lambda_{\min}{=}{-}L(L{-}1)/2$ as the largest and smallest eigenvalues of $G^{az}$. This results in $\mathcal{F}_{Q}{\leqslant}4n^{2}L^{2}(L-1)^{2}.$ Capturing this ultimate precision in estimating $h_{a}$ requires satisfying two conditions. First, $\alpha_{j}{\simeq}1$, that requires $\varepsilon{=}0$ and $h_{a}{\ll}\sqrt{2}\pi^{2}/L$. Second, the probe is to be initialized in the maximally entangled state as $\frac{1}{\sqrt{2}}\big{(}|{\uparrow,\cdots,\uparrow}\rangle_{a}{\otimes}|{\downarrow,\cdots,\downarrow}\rangle_{b}{\pm}|{\downarrow,\cdots,\downarrow}\rangle_{a}{\otimes}|{\uparrow,\cdots,\uparrow}\rangle_{b}\big{)}$. In the following, we show that, although initialized in the product state $|\psi(0)\rangle$, achieving this ultimate precision even for $\varepsilon{\neq}0$ is permitted in our DTC probe.
Fig. 2 (a) presents the QFI as a function of $h_{a}$ for various cycling periods when the periodic-field is in resonance with the binary quench, namely $T_{ac}{=}T$. Obviously, the QFI behaves differently in each phase. While in the DTC phase, one observes a plateau that uniformly increases with $n$, in the non-DTC region, non-trivial oscillations and suppression of the QFI are evident. The QFI clearly peaks at the same transition point denoted by $h_{a}^{\max}$ (dashed line) in Fig. 1 (b). The peaks become sharper as $n$ increases. Fig. 2 (b) depicts the dynamics of the QFI in three various regions, deep inside the DTC ($h_{a}{=}10^{-5}$), at the transition point ($h_{a}{=}h_{a}^{\max}$) and deep inside the non-DTC ($h_{a}{=}1$). The results show the square growth of the QFI, namely $\mathcal{F}_{Q}{\propto}n^{2}$ in the DTC phase. The unlimited quantum correlations at the transition point result in a faster growth of the QFI over early multiple cycles. However, at later times, one has $\mathcal{F}_{Q}{\propto}n^{2}$. In the Non-DTC region, the origin of QFI growth in later time is the contribution of several Floquet states that have considerable overlap with $|\psi(0)\rangle$ [111]. To extract the scaling behavior of the QFI, in Fig. 2 (c) we plot the obtained QFI as a function of $h_{a}$ after $n{=}10$ cycling period when the size of the chains varies. In both DTC and the transition point, the finite-size effect is apparent. By enlarging the chains, the peak of the QFI as the indicator of the phase transition skew toward smaller values of $h_{a}$ with a rate of $h_{a}^{\max}{\propto}L^{-1.0916}$, as shown in Fig. 2 (d). The result aligns with the analytical requirements of $h_{a}{\ll}\sqrt{2}\pi^{2}{/}L$ to achieve the ultimate precision. Although this result indicates degradation of the DTC order in the thermodynamic limit due to the presence of the periodic-field, in practice, it facilitates ultraprecise sensing of infinitesimal periodic-fields. In Fig. 2 (d), we show the values of the QFI deep inside the DTC phase ($h_{a}{=}10^{-5}$) and at the transition point ($h_{a}{=}h_{a}^{\max}$) when the length of the chains varies. The numerical results (markers) are well described by fitting the function $\mathcal{F}_{Q}{\propto}L^{\beta}$ with $\beta{=}4.0902$ and $\beta{=}2.5887$ for a sensor working inside the DTC order and at the transition point, respectively. This evidence indicates that leveraging a DTC system that is initialized in a product state $|\psi(0)\rangle$ results in a significant quantum enhancement in the sensing of the periodic-field.
To understand the origin of this quantum enhancement, one can focus on the energy spectrum of the period-doubling DTCs, namely the emergence of the ”Schrödinger cat” states [67, 70, 75, 111]. In the absence of the periodic-field, the spectrum $U_{n}$ in our DTC has a pair of eigenstates with strong overlap with $|\psi\rangle$ and $|\tilde{\psi}\rangle$ [111]. These eigenstates can be well approximated by the long-range correlated ”Schrödinger cat” states $|\Phi_{\pm}\rangle{\cong}\frac{1}{\sqrt{2}}(|\psi(0)\rangle{\pm}|\tilde{\psi}(0)\rangle)$, and the $\pi$ quasienergy gap between them causing the subharmonic response in dynamic form $|\psi(0)\rangle$ or $|\tilde{\psi}(0)\rangle$. For a small magnetic field, the spectrum of $U_{n}$ remains unchanged, and the sensing protocol can benefit from the freely generated maximally entangled states.
V Robustness analysis
Having elucidated the performance of the DTC probe for sensing a periodic-field, we now explore the viability of our probe under various types of defects. After relaxing the criteria of resonance in the periodic-field, we follow up by studying the effect of the crosstalk and perturbed initialization. We complete the study by analyzing the effect of the imperfect quench.
V.1 Off-resonance effect
Figure 3: (a) The imbalance as a function of $h_{a}$ for different values of the detuning $\delta f$ over different stroboscopic times. (b) The QFI versus the detuning $\delta f$ for various values of the $h_{a}$ and $n$. In both panels $L{=}7$. (c) The QFI in systems with different sizes as a function of $n$ when $h_{a}{=}10^{-2}$ and $\delta f{=}10^{-2}$. (d) The QFI as a function of $L$ when the sensor is in the DTC phase and at the transition point and $\delta f{=}10^{-2}$. The numerical results are well described by the fitting function $\mathcal{F}_{Q}{\propto}L^{\beta}$ with $\beta{=}4.6658$ and $\beta{=}2.2014$ for a sensor deep inside the DTC phase and at the transition point, respectively. In all the panels we set $\varepsilon{=}0.1$.
While far off-resonance periodic-field has a negligible impact on the dynamic of the system, slight deviation denoted by $f_{ac}/f{=}1{+}\delta f$ with $\delta f$ as the frequency offset may affect the lifetime of the DTC order and the performance of the probe. In this case, one obtains $\Theta_{n}{=}\frac{(-1)^{n}h_{a}}{\pi(1+\delta f)}\big{(}\cos(n\pi\delta f){+}\cos((n{-}1)\pi\delta f)\big{)}$ as the accumulated phase in each period cycle. In Fig. 3(a) we present the imbalance as a function of $h_{a}$ over multiple period cycles when $\delta f$ varies. The response of $\mathcal{I}(nT)$ in the DTC region reveals the robustness of the system to small offset values. Fig. 3 (b) plots the QFI as a function of $\delta f$ over multiple period cycles when the probe works inside the DTC order. Clearly, the maximum values for the QFI happen for a resonant periodic-field. By increasing $\delta f$, the unfavorable off-resonance correlations affect the performance of the probe. Note that $\Theta_{n}{\rightarrow}0$ whenever $n\delta f{\rightarrow}0.5$. This affects the growth of the accumulation phase and therefore the value of the QFI as can be seen in Fig. 3 (c) where we plot the dynamics of the QFI in systems of various lengths under a periodic-field with magnitude $h_{a}{=}10^{-2}$ and offset $\delta f{=}10^{-2}$. All the accumulated phase in the first time window, namely for $\delta fn{<}0.5$, will be erased in the second time window, namely for $\delta fn{>}0.5$. This introduces an optimal sensing time as $n^{*}{=}1/2\delta f$, giving the peaks of the QFI. Interestingly, sensing in $n{\leqslant}n^{*}$ kept the scaling behavior of the QFI intact, as evidenced by Fig. 3 (d), where we plot the obtained $\mathcal{F}_{Q}$ for different system sizes when the probe works inside the DTC and at the transition point with offset $\delta f{=}10^{-2}$. Clearly, the obtained scaling $\beta$ is in line with our results for the resonance case, indicating the attainability of the quantum-enhanced sensitivity.
So far, we have focused on the scenario where the periodic-field is detected exclusively by subsystem $a$. However, in practice, the close distance between chains $a$ and $b$ introduces crosstalk, causing $b$ also to sense a portion of the periodic-field. When initializing the subsystems in polarized states, but with opposite total magnetization in the $z$-direction, the crosstalk influences the accumulated phase and, consequently, the value of the QFI.
V.2 Cross-talk effect
Figure 4: (a) The QFI as a function of $h_{a}$ over various stereoscopic times, when $\eta{=}10^{-1}$. (b) The QFI as a function of $\eta$ when $n{=50}$. In both plots we set $L{=}7$ and $\varepsilon{=}0.1$.
To analyze the effect of the crosstalk, we replace the $H_{ac}(t)$ in Eq. (6) by
$H_{ac}(t)=B_{a}(t)G^{az}+\eta B_{a}(t)G^{bz},\quad\mathrm{with}\quad G^{\mu z}=\sum_{j=1}^{L}j\sigma^{\mu z}_{j},$ (12)
for $\mu{=}a,b$ in which $\eta{\in}(0,1)$. As evidenced by our results in Fig. 4 (a), where the QFI, as a function of $h_{a}$, has been plotted at different stroboscopic times when $\eta{=}10^{-1}$, the crosstalk can not affect the DTC order and the phase transition is driven by $h_{a}$. The distinct behavior of the QFI in both DTC and non-DTC regions, along with the enhanced peaks at the phase transition point, is evident. For completeness, in Fig. 4 (b) we present the QFI obtained as a function of $\eta$ when $h_{a}$ varies. The results are obtained after $n{=}50$ period cycling. Clearly, as the crosstalk parameter is increased, the values obtained for QFI decrease.
V.3 Initialization effect
Having analyzed the effect of off-resonance and crosstalk, we now explore the robustness of our DTC probe to imperfections in the initialization. Regardless of the presence or absence of the periodic-field, as a perfect quench with $\varepsilon{=}0$ preserves $z$-polarization in each chain, the DTC order with arbitrary initialization is infinitely long-lived even in systems with finite sizes. For $\varepsilon{>}0$, the configuration basis states of $H_{I}$ exhibit varying levels of robustness to imperfections, with some remaining resilient at higher values of $\varepsilon$. In the previous section, we showed that initializing the system in state $|\psi(0)\rangle{=}|{\uparrow,\cdots,\uparrow}\rangle_{a}{\otimes}|{\downarrow,\cdots,\downarrow}\rangle_{b}$ provides persistence period-doubling oscillations of $\langle\mathcal{I}(nT)\rangle$ even for $\varepsilon{\neq}0$. Furthermore, significant overlap of this initial state with the ”Schrödinger cat” state $|\phi_{\pm}\rangle{=}\frac{1}{\sqrt{2}}(|\psi(0)\rangle{\pm}|\tilde{\psi}(0)\rangle)$ in the spectrum of the DTC has been identified as the main source of quantum-enhanced sensing. This obliges us to analyze the effect of perturbation in the initialization of the probe. For this, we consider a product state as
$|\psi(0)\rangle$ $=\Pi_{j=1}^{L}|\varsigma_{j}\rangle_{a}\otimes|\varsigma_{j}\rangle_{b}$ (13)
$|\varsigma_{j}\rangle_{a}$ $={\rm cos(\alpha)}|\uparrow_{j}\rangle_{a}+{\rm sin(\alpha)}|\downarrow_{j}\rangle_{a}$ (14)
$|\varsigma_{j}\rangle_{b}$ $=-{\rm sin(\alpha)}|\uparrow_{j}\rangle_{b}+{\rm cos(\alpha)}|\downarrow_{j}\rangle_{b}$ (15)
with $\alpha{\in}[0,\pi/4]$, to avoid mixtures that yield $\langle\mathcal{I}(0)\rangle{=}0$. For the extreme values of $\alpha{=}0$, one can recover $|\psi(0)\rangle$. However, for other values of $\alpha$, a weighted mixture of different configurations is attainable. In Fig. 5 (a), we represent the normalized imbalance $\langle\mathcal{I}(nT)\rangle/\langle\mathcal{I}(0)\rangle$ over tens of period cycles when $\alpha$ varies. Clearly, the period-doubling oscillation of DTC remains unaffected by small imperfections during the initialization stage. To assess the performance of our DTC probe, in Fig. 5 (b), we plot the obtained QFI as a function of $h_{a}$ over different stroboscopic times when $\alpha{=}10^{-2}$. The results show that this level of perturbation in the initial state has no tangible impact on the performance of our DTC probe.
Figure 5: (a) The imbalance as a function of $n$ for different values of $\alpha$ when $h_{a}{=}10^{-2}$. (b) The QFI versus $h_{a}$ over different stroboscopic times when $\alpha{=}10^{-2}\pi$. In both panels $L{=}7$ and $\varepsilon{=}0.1$.
V.4 Imperfect quench effect
After explaining the sharp phase transition driven by $h_{a}$, we will now investigate the melting of the DTC by increasing the imperfection $\varepsilon$. In Fig. 6(a), we plot the imbalance $\mathcal{I}(nT)$ after $n{=}50$ period cycles as a function of $h_{a}$ and $\varepsilon$. The phase diagram is entirely characterized by the parameters $h_{a}$ and $\varepsilon$. Increasing $\varepsilon$ reduces the DTC phase imbalance, but can improve the performance of the probe. As evident in Fig. 6(b), where we plot the obtained QFI as a function of $\varepsilon$ deep inside the DTC phase and at the transition point, by increasing the imperfection in the quench up to some value, the QFI grows. This growth in the DTC phase is more tangible than at the transition point. During the DTC phase, the dynamics is limited to a subspace spanned by $|\Phi_{\pm}\rangle$. Increasing $\varepsilon$ by engaging a larger portion of the Hilbert space encodes more information about $h_{a}$ in the quantum states. This finding contrasts sharply with typical sensors, where imperfections diminish performance.
Figure 6: (a) The imbalance as a function of $h_{a}$ and $\varepsilon$ at $n{=}50$. (b) The QFI versus $\varepsilon$ for different values of the $h_{a}$ at $n{=}20$. In both panels $L{=}7$.
VI Experimental proposal
The methodology presented here is independent of the underlying physics, making it applicable to a broad spectrum of quantum and even classical systems. Here, we provide a proof-of-principle demonstration of our DTC sensor in ultracold atoms on an optical lattice, paving the way for realistic implementations.
Using the Jordan-Wigner transformation, one can easily map the Hamiltonian Eq. (4) to the spinless fermion model as
$H(t)=\begin{cases}H_{a}{+}H_{b}={-}J_{z}\sum_{j=1}^{L-1}\big{(}n^{a}_{j}n^{a}_{j+1}+n^{b}_{j}n^{b}_{j+1}\big{)}&0{<}t{\leqslant}t_{1}\\ H_{I}=J_{ab}\sum_{j=1}^{L}(c^{a\dagger}_{j}c^{b}_{j}+c^{a}_{j}c^{b\dagger}_{j})&t_{1}{<}t{\leqslant}t_{1}{+}t_{2},\end{cases}$ (16)
where $n^{\mu}_{j}{=}c^{\mu\dagger}_{j}c^{\mu}_{j}$ counts the particle numbers at site $j$ in chain $\mu{=}a,b$, and $c^{\mu\dagger}_{j}$ ($c^{\mu}_{j}$) is the fermionic creation (annihilation) operator. In this notation, the AC magnetic field can be rewritten as
$H_{ac}(t){=}B_{a}(t)\sum_{j=1}^{L}jn^{a}_{j}.$ (17)
Certain fermionic atoms, such as dysprosium (Dy) and erbium (Er), exhibit significant magnetic dipole moments. When cooled to ultracold temperatures and loaded into an optical lattice, these atoms can demonstrate behaviors governed by magnetic dipolar interactions. The magnetic dipole-dipole interactions, which are long-range and anisotropic, take the following form
$V_{jj^{{}^{\prime}}}=\frac{C_{dd}}{|r_{jj^{{}^{\prime}}}|^{3}}\Big{(}1-3(\hat{d}_{j}.\hat{r}_{jj^{{}^{\prime}}})(\hat{d}_{j^{{}^{\prime}}}.\hat{r}_{jj^{{}^{\prime}}})\Big{)}$ (18)
where $C_{dd}$ is the dipole-dipole interaction strength, $\hat{r}_{jj^{{}^{\prime}}}{=}\vec{r}_{j}-\vec{r}_{j^{{}^{\prime}}}/|\vec{r}_{j}-\vec{r}_{j^{{}^{\prime}}}|$, and $r_{jj^{{}^{\prime}}}{=}|\vec{r}_{j}-\vec{r}_{j^{{}^{\prime}}}|$. The unit vector $\hat{d}_{j}$ denotes the direction of the magnetic dipole of an atom at the site $j$, with coordinate $\vec{r}_{j}{=}x_{j}\hat{x}{+}y_{j}\hat{y}$, which is usually controlled by applying a uniform magnetic field. Implementing our DTC system requires an engineered optical potential to form a 1D lattice as well as a double-well potential with controllable barriers. During $t_{1}$, one can suppress the interaction between two chains $a$ and $b$ by ramping up the barrier of the double-well. Therefore, the dipolar interaction between the atoms with distance $x$ in each chain, with strength $C_{dd}/x{|j_{\mu}{-}j^{{}^{\prime}}_{\mu}|}$, generates $H_{a}{+}H_{b}$ in Eq. (4). Despite its long-range nature, its decay to the third power with distance likely does not impact the main results. To implement the tunneling between two chains $a$ and $b$, one needs to lower the barrier of the double-well, increase the optical potential to suppress the intra-chain interaction, and apply a uniform magnetic field as $\vec{B}_{2}{=}\sqrt{1/3}\hat{x}{+}\sqrt{2/3}\hat{z}$, that results in
$V_{j_{a}{\neq}j^{{}^{\prime}}_{a}}$ ${=}0,\quad V_{j_{b}{\neq}j^{{}^{\prime}}_{b}}{=}0,\quad V_{j_{a}{=}j_{b}}{=}\frac{C_{dd}}{y^{3}}$ (19)
$V_{j_{a}{\neq}j_{b}}$ ${=}\frac{C_{dd}}{r_{ab}^{3}}(1-\frac{|j_{a}-j_{b}|^{2}x^{2}}{r_{ab}^{2}}),$ (21)
where $r_{ab}{=}\sqrt{y^{2}+|j_{a}-j_{b}|^{2}x^{2}}$. Setting $y{\gg}x$, leads to $V_{j_{a}{=}j_{b}}{\rightarrow}0$ and $V_{j_{a}{\neq}j_{b}}{\rightarrow}0$ and allows tunneling for duration $t_{2}{\simeq}\hbar\pi y^{3}(1{-}\varepsilon){/}2C_{dd}$.
Imposing this probe on a periodic-field as $\vec{B}_{a}(nt){=}(-1)^{n}h_{a}\hat{y}$, while may not affect the intra-chain dipole-dipole interaction, results in
$V_{j_{a}{\neq}j^{{}^{\prime}}_{a}}$ ${=}0,\quad V_{j_{b}{\neq}j^{{}^{\prime}}_{b}}{=}0,\quad V_{j_{a}{=}j_{b}}{=}\frac{C_{dd}}{y^{3}}(1{-}\frac{x^{2}}{y^{2}}(1{+}\sqrt{3}(-1)^{n}j_{a}h_{a}\frac{y}{x}))$ (22)
$V_{j_{a}{\neq}j_{b}}$ ${=}\frac{C_{dd}}{r_{ab}^{3}}(1-\frac{(1+\sqrt{3}\frac{(-1)^{n}j_{a}h_{a}y}{|j_{a}-j_{b}|x})|j_{a}-j_{b}|^{2}x^{2}}{r_{ab}^{2}}).$ (23)
This result shows that, although, for infinitesimal $h$ and $y{\gg}x$, one can recover Eq. (19), larger values of $h$ affect the DTC nature through inducing site-dependent dipole-dipole interactions.
VII Conclusion
In this paper, we propose a DTC in a disorder-free system for sensing a gradient periodic-field with quantum-enhanced sensitivity. Our analysis results in $\mathcal{F}_{Q}{\leqslant}4n^{2}L^{2}(L{-}1)^{2}$ as the ultimate precision limit in the estimation of the gradient periodic-field. Our results show that a many-body probe initialized in a product state can achieve ultimate precision, namely $\mathcal{F}_{Q}{\propto}n^{2}L^{4}$, provided that it operates in the DTC phase. This enhancement is due to the resonance dynamics between two GHZ-type eigenstates of the system. Remarkably, this quantum enhancement shows robustness against different imperfections in the sensing protocol. In addition, we discuss the implementation of our probe in an optical lattice setup. Our results open avenues for harnessing inherent entanglement in interacting, non-equilibrium quantum many-body systems for enhanced metrology.
VIII Acknowledgment
AB acknowledges support from National Natural Science Foundation of China (Grants Nos. $12050410253$, $92065115$, and $12274059$) and the Ministry of Science and Technology of China (Grant No. QNJ2021167001202116700120211670012021167001L.